# Which Demographics do LLMs Default to During Annotation?

### v1

@@ 5 Conclusion and Future Work
With this paper, we presented the first analysis of the effect of socio-demographic prompting on tasks in the Popquorn data set. This corpus has been developed particularly for the study of the impact of demographic variables and therefore also offers itself for LLM-based studies.
@@
The findings are not overall consistent across models. While this could be considered a principled issue, it also presents itself with a substantial challenge: depending on the model that is used by an end-user, the impact of socio-demographic information varies, and different models default to different demographic information.
We therefore propose that future work focuses on the development of reliable instruments to measure the influence of implicitly or explicitly provided information about the user. We presume that making such information explicit and ensuring that it is interpreted in a reliable manner by models is an appropriate approach to mitigate issues with unknown biases that are not transparent to users or automatic processing pipelines.
@@
Our work has the goal to make challenges transparent which the use of large language models has. A considerable limitation from an ethical perspective is that the variables we consider is not relevant across cultures in the world. Particularly the “race” variable is heavily biased towards the US American society. Other variables might be relevant in other cultures in the world that we are not aware of. We therefore suggest to conduct future studies that consider more open sets of variables that describe the users of language model-based systems.
@@ A.1 Full Results for LLMs Mimicing Annotators
BIC (Bayesian Information Criterion) is a model fit statistic useful for model comparison, though its value is not meaningful in isolation. It is included to allow for calculations of BIC values for both base and limited models. The ICC (Intraclass Correlation Coefficient) indicates the extent to which the clustering (by instances and annotators) influences the outcomes (distances between human and LLM ratings). It represents the ratio of total variance in the dependent variable attributed to variance between cluster means, as opposed to variance within clusters. A value of 0.3 or higher suggests that the inclusion of random intercepts is necessary to account for clustering. If the ICC were very low (less than 0.1), a simpler model could be justified.
Finally, RMSE (Root Mean Square Error) represents the standard deviation of the residuals; however, it provides limited utility in isolation.
@@ A.3 Results for Placebo Prompting
Table 8 presents the results for placebo prompting. It includes sample sizes and mean distance scores ($\Delta_{\mu}$) for predictions generated using placebo prompts compared to those produced with no-info prompts for the models GPT-4o and Claude across two rating tasks. The results show that placebo prompting (P prompts) does not yield any notable differences relative to no-info prompting (N prompts) for specific attribute values. Overall, the scores remain stable across these comparisons. This consistency suggests that introducing non-relevant information does not change the models’ performance systematically. It reinforces our conclusion that the notable changes in model behavior are primarily driven by demographic prompting rather than by placebo or non-relevant prompts.

### v2

@@ 5 Conclusion and Future Work
With this paper, we present the first analysis of the effect of socio-demographic prompting on tasks in the Popquorn data set. This corpus has been developed particularly for the study of the impact of demographic variables and therefore also offers itself for LLM-based studies.
@@
Not all of our findings are overall consistent across models. While this could be considered an issue that is inherent to LLMs, it also presents itself with a substantial challenge: depending on the model that is used by an end-user, the impact of socio-demographic information varies, and different models default to different demographic information.
We therefore propose that future work should focus on the development of reliable instruments to measure the influence of both implicitly and explicitly provided demographic information about the user. We presume that making such information explicit and ensuring it is reliably interpreted by models is crucial not only to mitigate issues related to unknown biases that are not transparent to users or automatic processing pipelines, but also to create models that accurately replicate the diversity and variance across different individuals.
@@
Our work has the goal to make challenges transparent which the use of large language models has. A considerable limitation from an ethical perspective is that the variables we consider are not relevant across cultures in the world. The response options of the “race” variable, in particular, are specific to US American society. Other variables might be relevant in other cultures in the world that we are not aware of. We therefore suggest to conduct future studies that consider more open sets of variables that describe the users of language model-based systems.
Beyond the limitations of variable selection, we must also reflect on the implications of our findings for issues of inequality in the context of large language models. If models effectively mimic particular demographic groups, this can lead to the privileging of the viewpoints and experiences of those groups, potentially marginalizing individuals who are not adequately represented. Furthermore, if the responses are influenced by the inclusion of certain demographic information over others, it suggests that some demographic categories are represented as more “normal” or typical than others. This pattern mirrors historical inequities that persist in new technologies, revealing an unsettling continuity in the ways that cultural biases may be perpetuated.
@@ A.1 Full Results for LLMs Mimicing Annotators
The ICC (Intraclass Correlation Coefficient) indicates the extent to which the clustering (by instances and annotators) influences the outcomes (distances between human and LLM ratings). It represents the ratio of total variance in the dependent variable attributed to variance between cluster means, as opposed to variance within clusters. A value of 0.3 or higher suggests that the inclusion of random intercepts is necessary to account for clustering. If the ICC were very low (less than 0.1), a simpler model could be justified.
@@ A.3 Results for Placebo Prompting
Table 8 presents the results for placebo prompting. It includes sample sizes and mean distance scores ($\Delta_{\mu}$) for predictions generated using placebo prompts compared to those produced with no-info prompts for the models GPT-4o and Claude across two rating tasks. The results show that placebo prompting (P prompts) does not yield any notable differences relative to no-info prompting (N prompts) for specific attribute values. Overall, the scores remain stable across these comparisons.